# Detecting Spike Wave Discharges (SWD) using 1-dimensional Residual UNet

### v1

@@ Results
Because AugUNet1D performs well when labeling SWD, we believe that others attempting to automatically label events in EEG may find it useful. While our training dataset was quite large (22,637 events), others may have fewer examples for training. To estimate the quantity of training data needed to achieve adequate performance, we investigated the data efficiency of our proposed AugUNet1D model by training on varying fractions of the available training data (5%, 10%, 25%, 50%, 75%, and 90%) and evaluating performance on the same held-out test set. The results demonstrate remarkable data efficiency while revealing important insights about the model’s learning dynamics and the role of training data quantity in cross-subject generalization.
The results are provided in Table 4.
@@
Even with only 5% of the training data, the model achieved surprisingly strong performance with an F1-score of 0.8192, maintaining high precision (0.9350) and reasonable recall (0.7289). This result indicates that the combination of our residual U-Net architecture and comprehensive augmentation strategy enables effective learning from limited labeled data which can be a critical advantage for seizure detection applications where obtaining large annotated datasets is excessively laborious, expensive, or time-consuming. Performance improvements were most pronounced when increasing training data from 10% to 25%, with F1-score jumping to 0.8543. This substantial gain indicates that while the model can learn basic seizure patterns from minimal data, exposure to a more diverse set of training examples significantly enhances generalization.
Beyond 50% of training data, performance improvements plateaued, with 75% and 90% of data yielding F1-scores of 0.8524 and 0.8618, respectively.
Examining the precision-recall dynamics across different training data fractions reveals interesting patterns. Precision remained remarkably stable across all conditions (0.9184-0.9353), while recall showed the primary improvements as training data increased (0.7166 at 10% to 0.8007 at 90%). This pattern indicates that the model maintains consistent specificity regardless of training set size, but requires more diverse training examples to improve its sensitivity to varied seizure manifestations.
These findings have important practical implications for deployment of seizure detection systems. The strong performance achieved with only 25-50% of training data suggests that effective models can be developed without requiring exhaustive data collection from large numbers of subjects. However, the plateau beyond 50% also indicates that simply collecting more data from the same subject population may not substantially improve cross-subject performance—instead, improvements may require either architectural innovations, more sophisticated augmentation strategies, or inclusion of training data from more diverse subject populations with varied seizure characteristics.
@@
We conducted a hyperparameter study to determine the optimal probability (p) for applying amplitude scaling augmentation during training, evaluating probabilities ranging from 0.1 to 0.5. The results reveal a critical trade-off between precision and recall that is directly influenced by the augmentation probability, with substantial implications for model robustness and clinical utility.
@@
A notable exception emerged at p=0.2, which achieved competitive performance with an F1-score of 0.8577 through a more balanced precision-recall profile (0.9239 precision, 0.8003 recall). This suggests that a moderate augmentation frequency provides some improvement in recall while maintaining strong precision. However, increasing the augmentation probability to p=0.5 produced the optimal performance with an F1-score of 0.884, representing the best overall detection capability. This configuration achieved a substantial gain in recall to 0.8618 while maintaining strong precision at 0.9092, successfully detecting approximately 86% of seizure events with 91% of positive predictions being correct.
@@ Discussion
Reliably and accurately detecting events of interest in EEG signals is a common problem. Manually scanning through EEG traces and demarcating relevant intervals is time intensive and can be affected by extraneous factors like scorer experience, bias, and attention level. Attempts at automated approaches to these problems often involve using features in the time and frequency domains of the EEG (e.g. spectral profiles, changes in powers of specific frequency bands, waveform shape)[30, 31, 32, 23], machine learning [5, 33], or a combination of both [10, 34]. However, for the problem of detecting frequent, spontaneous SWD in long continuous recordings from C3H/HeJ mice, we found available methods inadequate. Inspired by U-Nets used in image segmentation and classification[16, 35, 36], we applied the same neural network architecture to the problem of SWD detection. The result is AugUnet1D, a U-Net neural network architecture for 1-dimensional time series data that uses data augmentation strategies during training to optimize detection of SWD in EEG.
@@
After training the network with 961 hours of recording, including 22,637 SWDs labeled manually by experienced researchers, AugUNet1D was tested on 10 EEG recordings from 10 different mice not included in the original training dataset. It achieved an overall average F1 score of 0.90 with a recall of 0.89 and precision of 0.90 (2). By comparison with all other models, it is clear that the combination of AugUNet1D’s encoder-decoder architecture with skip connections, residual learning for gradient stability, and data augmentation are beneficial for achieving state-of-the-art performance on cross-subject seizure detection. The U-Net architecture implemented in AugUNet1D permits the network to create hierarchical temporal representations, which explain much of its success on this data labeling task. Traditional machine learning approaches like Decision Tree and Logistic Regression performed far worse than AugUNet1D, likely due to their limited capacity to learn complex hierarchical representations from raw EEG signals. Recurrent architectures including LSTM, biLSTM, and ConvLSTM also failed despite their theoretical suitability for sequential data. The inability of these models to produce any meaningful predictions indicates that temporal modeling alone, without spatial feature extraction through convolution, is insufficient for this task. Other architectures, like EEGNet for example, likely fail because they were designed for multi-channel spatial-spectral feature extraction which differs fundamentally from the dense temporal segmentation required for SWD detection. Finally, a recently published time- and frequency-based method, Twin Peaks, was also inferior to AugUNet1D and has lower temporal resolution. AugUNet then represents the best overall method for SWD detection in our dataset.
@@
In addition to performing better overall, one especially important advantage of AugUNet1D is that it can very precisely segment SWDs. It marks the starts and ends of detected events very close to where trained researchers draw event boundaries (see Figure 5 for examples). Time- and frequency-based detection methods rely on changes in the power spectrum, which has poor temporal resolution as a consequence of computing a Fourier Transform over a finite time window. Therefore, the time resolution of the detection signal is only a fraction of the time resolution of the original EEG data acquired. Similarly, machine learning applied to chunked data [37] or pre-detected candidate events[10] also presents the same segmentation problem: the beginning and end of an event lack temporal resolution. AugUNet1D avoids this problem altogether by reading in all the raw data and treating every sample as a discrete point to label rather than grouping them into blocks with homogeneous compositions.
@@
Another notable feature of AugUNet1D is its robust performance during time periods when other SWD detection methods fail. Time- and frequency-based methods tend to falsely detect events during sleep when the overall EEG amplitude increases. Therefore, methods that rely on changes in EEG amplitude to detect SWDs also tend to detect events during sleep (Figure 6B). Transient, non-pathological EEG oscillations during sleep can increase power in the SWD frequency band (5-7Hz for C3H/HeJ mice ) substantially and are often erroneously detected by thresholding in that frequency band. Overall, false detections during sleep pose a serious problem for time- and frequency-based detection methods. SWDs can happen during sleep[38, 39] so excluding sleep altogether is not an adequate solution. AugUNet1D presents a superior alternative by learning the SWD pattern itself rather than its spectral composition or simplified waveform features which can be shared by non-pathological EEG events.
@@
In conclusion, AugUNet1D is more effective for automatically identifying SWDs in our dataset than all other 16 methods attempted here. It outperforms all other traditional machine learning methods, neural networks, and the Twin Peaks algorithm. Furthermore, AugUNet1D is robust, performing well even in conditions where other methods fail, like during slow wave sleep. We anticipate that AugUNet1D will also be effective in detecting other EEG events more generally and that will be valuable for experimenters and clinicians alike; data presented here show that trained researchers spend almost 1 hour manually labeling 9 hours of continuous EEG data. AugUNet1D can perform this task in a fraction of the time and doesn’t require users to actively part during processing.

### v2

@@ Results
Because AugUNet1D performs well when labeling SWDs, we were motivated to evaluate model perforamnce when trained with fewer events (our training dataset was quite large: 22,637 SWDs, 173,160 20-second segments). To estimate the quantity of training data needed to achieve adequate performance, we investigated the data efficiency of our proposed AugUNet1D model by training on varying fractions of the available training data (5%, 10%, 25%, 50%, 75%, and 90%), and then evaluating performance on the same held-out test set. Our results (Table 4) demonstrate remarkable data efficiency while revealing important insights about the model’s learning dynamics and the role of training data quantity in cross-subject generalization.
@@
By limiting the training dataset to only 5%, the model achieved surprisingly strong performance with an F1-score of 0.8192, maintaining high precision (0.9350) and reasonable recall (0.7289). This result indicates that the combination of our residual U-Net architecture and comprehensive augmentation strategy enables effective learning from limited labeled data, a potentially critical advantage for seizure detection applications where obtaining large, annotated datasets is excessively laborious, expensive, or time-consuming. Performance improvements were most pronounced when increasing training data from 10% to 25%, with the F1-score increasing to 0.8543. This substantial gain indicates that while the model can learn basic seizure patterns from minimal data, exposure to a more diverse set of training examples significantly enhances generalization.
Beyond 50% of training data, performance improvements plateaued, with 75% and 90% of data yielding F1-scores of 0.8524 and 0.8618, respectively. Notably, evaluating the precision-recall dynamics across different training data fractions revealed interesting patterns. Precision remained remarkably stable across all conditions (0.9184-0.9353), whereas recall showed primary improvements as training data increased (0.7166 at 10% to 0.8007 at 90%). This pattern indicates that the model maintains consistent specificity regardless of training set size but requires more diverse training examples to improve its sensitivity to varied seizure manifestations.
These findings have important practical implications for the deployment of seizure detection systems. The strong performance achieved with only 25-50% of training data suggests that effective models can be developed without requiring exhaustive data collection from large numbers of subjects. However, the plateau beyond 50% also indicates that simply collecting more data from the same subject population may not substantially improve cross-subject performance. Instead, improvements may require either architectural innovations, more sophisticated augmentation strategies, or the inclusion of training data from more diverse subject populations with varied seizure characteristics.
@@
Finally, we conducted a hyperparameter study to resolve the optimal probability (p) for applying amplitude scaling augmentation during training, evaluating probabilities ranging from 0.1 to 0.5. Our results (5) revealed a critical trade-off between precision and recall that is directly influenced by the augmentation probability, with substantial implications for model robustness and clinical utility.
@@
A notable exception emerged at p=0.2, which achieved competitive performance with an F1-score of 0.8577 through a more balanced precision-recall profile (0.9239 precision, 0.8003 recall). This observation suggests that a moderate augmentation frequency provides some improvement in recall while maintaining strong precision. However, increasing the augmentation probability to p=0.5 produced the optimal performance with an F1-score of 0.884, representing the best overall detection capability. This configuration achieved a substantial gain in recall to 0.8618 while maintaining strong precision at 0.9092, successfully detecting approximately 86% of seizure events with 91% of positive predictions being correct.
@@ Discussion
Reliably and accurately detecting events of interest in EEG signals is a common problem. Manually scanning through EEG traces and demarcating relevant intervals is time intensive and can be affected by extraneous factors like scorer experience, bias, and attention level. Attempts to automatically detect salient EEG events often involve using features in the time and frequency domains of the EEG (e.g. spectral profiles, changes in powers of specific frequency bands, waveform shape)[30, 31, 32, 23], machine learning approaches [5, 33], or a combination of both [10, 34]. However, when detecting frequent, spontaneous SWDs observed during long continuous recordings from C3H/HeJ mice, we found available methods inadequate. Inspired by U-Nets used in image segmentation and classification[16, 35, 36], we applied the same neural network architecture to the problem of SWD detection. The result is AugUnet1D, a U-Net neural network architecture for 1-dimensional time series data that uses data augmentation strategies during training to optimize SWD detection.
@@
After training the network with 961 hours of EEG recording, including 22,637 SWDs labeled manually by experienced researchers, AugUNet1D was tested on 10 EEG recordings from 10 different mice not included in the original training dataset. It achieved an overall average F1 score of 0.90 with a recall of 0.89 and precision of 0.90 (2). By comparison with all other models, it is clear that the combination of AugUNet1D’s encoder-decoder architecture with skip connections, residual learning for gradient stability, and data augmentation significantly improve performance on cross-subject seizure detection. The U-Net architecture implemented in AugUNet1D permits the network to create hierarchical temporal representations, which explain much of its success on this data labeling task. Traditional machine learning approaches like Decision Tree and Logistic Regression performed far worse than AugUNet1D, likely due to their limited capacity to learn complex hierarchical representations from raw EEG signals. Recurrent architectures including LSTM, biLSTM, and ConvLSTM also failed despite their theoretical suitability for sequential data. The inability of these models to produce any meaningful predictions indicates that temporal modeling alone, without spatial feature extraction through convolution, is insufficient for this task. Other architectures (e.g., EEGNet) likely fail because they were designed for multi-channel spatial-spectral feature extraction which differs fundamentally from the dense temporal segmentation required for SWD detection. Finally, a recently published time- and frequency-based method, Twin Peaks, was also inferior to AugUNet1D and has lower temporal resolution. AugUNet then represents the best overall method for SWD detection for our dataset.
@@
In addition to performing better overall, one especially important advantage of AugUNet1D is that it can very precisely segment SWDs. It marks the starts and ends of detected events very close to where trained researchers draw event boundaries (see Figure 8 for examples). Time- and frequency-based detection methods rely on changes in the power spectrum, which has poor temporal resolution as a consequence of computing a Fourier Transform over a finite time window. Therefore, the time resolution of the detection signal is only a fraction of the time resolution of the original EEG data acquired. Similarly, machine learning applied to chunked data [37] or pre-detected candidate events[10] also presents the same segmentation problem: the beginning and end of an event lack temporal resolution. AugUNet1D avoids this problem altogether by reading the entire, continuous data and treating every sample as a discrete point to label rather than grouping samples into blocks with homogeneous compositions.
@@
Another notable feature of AugUNet1D is its robust performance during periods when other SWD detection methods fail. Time- and frequency-based methods tend to falsely detect events during sleep when the overall EEG amplitude increases. Therefore, methods that rely on changes in EEG amplitude to detect SWDs also tend to detect events during sleep (Figure 8C). Transient, non-pathological EEG oscillations during sleep can increase power in the SWD frequency band (5-7Hz for C3H/HeJ mice ) substantially and are often erroneously detected by thresholding in that frequency band. Overall, false detections during sleep pose a serious problem for time- and frequency-based detection methods. SWDs can happen during sleep[38, 39] so excluding sleep altogether is not an adequate solution. AugUNet1D presents a superior alternative by learning the SWD pattern itself rather than its spectral composition or simplified waveform features which can be shared by non-pathological EEG events.
@@
In conclusion, AugUNet1D is more effective for automatically identifying SWDs in our dataset than all other methods tested here. It outperforms all other traditional machine learning methods, neural networks, and the Twin Peaks algorithm. Furthermore, AugUNet1D is robust, performing well even in conditions where other methods fail (e.g., during slow-wave sleep). We also anticipate that AugUNet1D can be applied more generally to detect a range of EEG signals, including non-SWD EEG events and that AugUNet1D will be valuable for experimenters and clinicians alike; the data presented herein show that trained researchers require nearly one hour to manually label every nine hours of continuous EEG data. AugUNet1D can perform this task much more efficiently and does not require users to actively process data.
@@
Funding
This work was supported by the National Institutes of Health under grants R01 NS126594, R01 NS131670, and T32 DK007646.